2,The return on shares of Valley Transporter is predicted under the following various economic conditions:
Recession -0.15
Normal +0.07
Boom +0.18
If each economy state has the same probability of occurring, what is the variance of the stock?

3,The return on shares of the Orange Company are predicted under the following states of nature. The states of nature are all equally likely, and because there are a total of three states, each state has a 33.333% chance of occurring.

Recession -0.13
Normal +0.04
Boom +0.24

What is the standard deviation of Orange?

* Place your answer in decimal form, for example as say .0675 and not 6.75.
Answer #1

Answer to Question 2:

Expected Return = (1/3) * (-0.15) + (1/3) * 0.07 + (1/3) * 0.18
Expected Return = 0.0333

Variance = (1/3) * (-0.15 - 0.0333)^2 + (1/3) * (0.07 - 0.0333)^2 + (1/3) * (0.18 - 0.0333)^2
Variance = 0.01882

Answer to Question 3:

Expected Return = (1/3) * (-0.13) + (1/3) * 0.04 + (1/3) * 0.24
Expected Return = 0.05

Variance = (1/3) * (-0.13 - 0.05)^2 + (1/3) * (0.04 - 0.05)^2 + (1/3) * (0.24 - 0.05)^2
Variance = 0.02287

Standard Deviation = (0.02287)^(1/2)
Standard Deviation = 0.1512